What is the solution of the eqation 1/3(x+2)=2/3x+4

1
Expert's answer
2022-03-19T02:38:00-0400

1/3(x+2) = 2/3x+4

1/ (3x+2) = 2/ (3x+4)

Cross multiply

2(3x+2) = 3x+4

6x +12 = 3x +4

6x-3x = 4-12

3x = -8

Therefore, x = -8/3